OpenStudy (anonymous):

The formula for degrees Celsius is: C = 9 over 5 (F – 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 6 degrees Celsius is.

OpenStudy (anonymous):

Part 2: 6 degrees Celcius is 42.8 degrees Fahrenheit.

OpenStudy (anonymous):

C=5/9(F-32) 9/5*C=(F-32) 9/5*C+32=F F=C*9/5+32

OpenStudy (anonymous):

was trying to understand all those but. im super lost

OpenStudy (nincompoop):